# ActiView: Evaluating Active Perception Ability for Multimodal Large Language Models

### v1

@@ Appendix G Case Study of Providing Human-annotated Clues
Appendix H Limitations and Future Topics
In this study, we utilize a specific form of Visual Question Answering (VQA) to assess active perception abilities of models. While this form of VQA presents significant challenges for current multimodal language models (MLLMs), it does not encompass all aspects of active perception. For instance, it temporarily overlooks factors such as perspective distortion, multi-sensor integration, and the incorporation of more dynamic or interactive environments. Moreover, this study solely evaluates the inherent capabilities of the MLLMs. Techniques like tool learning and multi-agent collaboration could potentially enhance active perception performance based on existing MLLMs, making these areas worthy for future exploration and improvement. Also, the free-form evaluation of active perception can be included, as some agents possess the ability to freely crop the image and move their perceptual fields around upon external instructions and interactions. Given that these limitations exceed the scope of a single paper, we will leave them as future work.
Appendix I Prompt Template
@@
I.1 Templates for General Question Answering
@@
I.2 Templates for Zooming Evaluation
@@
I.3 Templates for Shifting Evaluation
@@
I.4 Templates for Mixed Evaluation
@@
An Example Prompt for Operation Determination (1) You will be presented with a full image <image> and a corresponding question to answer. The image is split in to <num_splits> equal parts, numbered from 1 to <num_splits>, where <description_of_splits>. You can check for detailed visual information via zooming operation that zoom in to your selected part or parts iwth from the above numbers. Response with the the numbers of parts you wish to zoom in, or response with ‘‘none’’ if you don’t need to can check for details.
@@
Appendix J Attempts of Automatic Data Generation
@@
J.1 Automatic Data Generation Process
@@
J.2 Cases of Unsuccessful Generations of GPT-4V
@@
J.3 Prompt Templates Used for Automatic Generation
@@
Heuristic prompt used for automatic annotation in Step 1 The clues for marking information in several bboxes in this picture are: {clues} Based on several bboxes and corresponding clues, please design a question that requires the model to synthesize the information in these bboxes (at least two, and can only be answered based on the information in the bboxes and the clues corresponding to the annotated information). You only need to ask the question, and there is no need to repeat the clue again. Note that the existence of bbox (including its ID information) cannot be mentioned in the question. Questions and reasoning should be based on objective facts as much as possible instead of subjective guessing.But at the same time, you should also avoid grounding questions and questions that can be answered without pictures (including questions like what someone in the picture is doing) Next, mark me the corresponding information in the following format: 1. ‘‘question’’ (str) 2. ‘‘options’’ (list) 3. ‘‘abilities’’ (list): choose from ‘‘zoom in’’, ‘‘zoom out’’, ‘‘shifting’’ (it is mentioned in the analysis and is not mentioned at the beginning) 4. ‘‘answer’’ (int, index of option) 5. ‘‘order’’: the order in which the pictures cut out of the bbox and the entire picture are displayed (the list is given in the order of reasoning, all of which are ints, representing the id corresponding to the bbox on the picture, if it is a complete picture, it is 0) 6. “groundtruth”: Give the reasons and complete reasoning process for answering the question 7. ‘‘number_of_operations’’: For example, first zoom in and then move the angle of view, it is two operations You must give me the answer in the following json-string format(not code block) and dont́ say anything else: {{ ‘‘question’’: question(str), ‘‘options’’: options(list), ‘‘abilities’’: ablities(list), ‘‘answer’’: answer_index(int), ‘‘order’’: order(list), ‘‘groundtruth’’: groundtruth(str), ‘‘number_of_operations’’: number of the operations(int)
@@
Scoring prompt used for automatic annotation in Step 1 We want to design a question about the picture to test the active perception ability of the respondent. Here are the requirements: You will be provided with an image and information of bboxes in it. You should design a question that requires the respondent to synthesize the information in these bboxes. While designing the questions, you must follow these rules: - The question should be based on the information in the given bboxes. - The question requires the respondent to obtain information from the field of view of these bboxes as a basis, identify irrelevant information on the picture, and move the field of view of different bboxes to obtain more information before answering the question. - Differences between options should be distinct. And options must not be conflict to each other. - There should be one and only one correct answer among all options. - The evidence or clues for answering the question must be visible in the image. Also, you should realize the following conditions: - The answers must not require the respondent guess subjectively. - You cannot generate questions require simple object grounding, e.g., what is the object in a certain region, what is the color of an object, etc. - The existence of bbox and visual clues (including their ID information) cannot be mentioned in the question nor in the options. You should score the annotation through the rules given above. Here are the predefined levels for scoring, where level D is the worst and level A is the best: - Level D: no reasonable questions can be generated for the given image by strictly following our rules. - Level C: the question contains subjective guesses and judgments, rather than strictly following the rules(e.g. infer the location from the architectural style/image style rather than some grounding signs and texts etc.) - Level B: the question can be answered via simple captioning of the pictures(like using ViT or OCR to caption the picture and ask the language model to answer the question with out the picture), or can be answered via pure common sense reasoning. - Level A: the question is cleverly designed and is completely based on the information in the picture. It requires the respondents to visit different parts marked on the image for comprehensive reasoning, which fully complies with the above marking rules. Remenber, if any subjective guess seems to appear, or anything that requires inferring from knowledge outside the image, or anything that does not follow our rule strictly (including asking for some weired questions etc.), do not hesitate to assign a low level. Here’s the annotation information of the given picture: {annotation} You must give me the answer in the following json-string format(not code block) and don’t say anything else: {{ ‘‘score’’: string, choose from ‘‘A’’, ‘‘B’’,‘‘C’’, ‘‘D’’, ‘‘reason’’: string, explain why you give this score
@@
3. ‘‘abilities’’ (list): choose from ‘‘zoom in’’, ‘‘zoom out’’, ‘‘shift’’ 4. ‘‘answer’’ (int, index of correction option, starting from 0) 5. “groundtruth”: Give the reasons and complete reasoning process for answering the question 6. ‘‘operations’’: For example, first zoom in to a region and then moving to a different region, counted as two operations === Here are some bounding boxes and their type for you to refer to: {boxes} The items in these bounding boxes are all {type} The questions you ask must be about the information within the bounding boxes and strictly meet the requirements and question types given to you above. === If it is impossible to come up with required questions, you should response with ‘‘question’’: (str)‘‘None’’ in json-string format(not code block). Otherwise, you must generate response in the following json-string format(not code block) and dont́ say anything else: {{ ‘‘question’’: question(str), ‘‘options’’: options(list), ‘‘abilities’’: ablities(list), ‘‘answer’’: answer_index(int), ‘‘order’’: order(list), ‘‘groundtruth’’: groundtruth(str), ‘‘operations’’: number of the operations(int), }} === Please generate response for the given image that **strictly follow** the above requirments:

### v2

@@ I.2 Cases of Giving Human-annotated Clues
Appendix J Prompt Template
@@
J.1 Templates for General Question Answering
@@
J.2 Templates for Zooming Evaluation
@@
J.3 Templates for Shifting Evaluation
@@
J.4 Templates for Mixed Evaluation
@@
An Example Prompt for Operation Determination (1) You will be presented with a full image <image> and a corresponding question to answer. The image is split in to <num_splits> equal parts, numbered from 1 to <num_splits>, where <description_of_splits>. You can check for detailed visual information via zooming operation that zoom in to your selected part or parts with from the above numbers. Response with the the numbers of parts you wish to zoom in, or response with ‘‘none’’ if you don’t need to can check for details.
@@
Appendix K Attempts of Automatic Data Generation
@@
K.1 Automatic Data Generation Process
@@
K.2 Cases of Unsuccessful Generations of GPT-4V
@@
K.3 Prompt Templates Used for Automatic Generation
@@
Heuristic prompt used for automatic annotation in Step 1 The clues for marking information in several bboxes in this picture are: {clues} Based on several bboxes and corresponding clues, please design a question that requires the model to synthesize the information in these bboxes (at least two, and can only be answered based on the information in the bboxes and the clues corresponding to the annotated information). You only need to ask the question, and there is no need to repeat the clue again. Note that the existence of bbox (including its ID information) cannot be mentioned in the question. Questions and reasoning should be based on objective facts as much as possible instead of subjective guessing.But at the same time, you should also avoid grounding questions and questions that can be answered without pictures (including questions like what someone in the picture is doing) Next, mark me the corresponding information in the following format: 1. ‘‘question’’ (str) 2. ‘‘options’’ (list) 3. ‘‘abilities’’ (list): choose from ‘‘zoom in’’, ‘‘zoom out’’, ‘‘shifting’’ (it is mentioned in the analysis and is not mentioned at the beginning) 4. ‘‘answer’’ (int, index of option) 5. ‘‘order’’: the order in which the pictures cut out of the bbox and the entire picture are displayed (the list is given in the order of reasoning, all of which are ints, representing the id corresponding to the bbox on the picture, if it is a complete picture, it is 0)
Heuristic prompt used for automatic annotation in Step 1 (Continue) 6. “groundtruth”: Give the reasons and complete reasoning process for answering the question 7. ‘‘number_of_operations’’: For example, first zoom in and then move the angle of view, it is two operations You must give me the answer in the following json-string format(not code block) and dont́ say anything else: {{ ‘‘question’’: question(str), ‘‘options’’: options(list), ‘‘abilities’’: ablities(list), ‘‘answer’’: answer_index(int), ‘‘order’’: order(list), ‘‘groundtruth’’: groundtruth(str), ‘‘number_of_operations’’: number of the operations(int)
@@
Scoring prompt used for automatic annotation in Step 1 We want to design a question about the picture to test the active perception ability of the respondent. Here are the requirements: You will be provided with an image and information of bboxes in it. You should design a question that requires the respondent to synthesize the information in these bboxes. While designing the questions, you must follow these rules: - The question should be based on the information in the given bboxes. - The question requires the respondent to obtain information from the field of view of these bboxes as a basis, identify irrelevant information on the picture, and move the field of view of different bboxes to obtain more information before answering the question. - Differences between options should be distinct. And options must not be conflict to each other. - There should be one and only one correct answer among all options. - The evidence or clues for answering the question must be visible in the image. Also, you should realize the following conditions: - The answers must not require the respondent guess subjectively. - You cannot generate questions require simple object grounding, e.g., what is the object in a certain region, what is the color of an object, etc. - The existence of bbox and visual clues (including their ID information) cannot be mentioned in the question nor in the options. You should score the annotation through the rules given above. Here are the predefined levels for scoring, where level D is the worst and level A is the best: - Level D: no reasonable questions can be generated for the given image by strictly following our rules. - Level C: the question contains subjective guesses and judgments, rather than strictly following the rules(e.g. infer the location from the architectural style/image style rather than some grounding signs and texts etc.) - Level B: the question can be answered via simple captioning of the pictures(like using ViT or OCR to caption the picture and ask the language model to answer the question with out the picture), or can be answered via pure common sense reasoning. - Level A: the question is cleverly designed and is completely based on the information in the picture. It requires the respondents to visit different parts marked on the image for comprehensive reasoning, which fully complies with the above marking rules.
Scoring prompt used for automatic annotation in Step 1 (Continue) Remenber, if any subjective guess seems to appear, or anything that requires inferring from knowledge outside the image, or anything that does not follow our rule strictly (including asking for some weired questions etc.), do not hesitate to assign a low level. Here’s the annotation information of the given picture: {annotation} You must give me the answer in the following json-string format(not code block) and don’t say anything else: {{ ‘‘score’’: string, choose from ‘‘A’’, ‘‘B’’,‘‘C’’, ‘‘D’’, ‘‘reason’’: string, explain why you give this score
@@
3. ‘‘abilities’’ (list): choose from ‘‘zoom in’’, ‘‘zoom out’’, ‘‘shift’’ 4. ‘‘answer’’ (int, index of correction option, starting from 0) 5. “groundtruth”: Give the reasons and complete reasoning process for answering the question 6. ‘‘operations’’: For example, first zoom in to a region and then moving to a different region, counted as two operations $=$
Prompt used for automatic annotation in Step 2 (Continue) Here are some bounding boxes and their type for you to refer to: {boxes} The items in these bounding boxes are all {type} The questions you ask must be about the information within the bounding boxes and strictly meet the requirements and question types given to you above. === If it is impossible to come up with required questions, you should response with ‘‘question’’: (str)‘‘None’’ in json-string format(not code block). Otherwise, you must generate response in the following json-string format(not code block) and dont́ say anything else: {{ ‘‘question’’: question(str), ‘‘options’’: options(list), ‘‘abilities’’: ablities(list), ‘‘answer’’: answer_index(int), ‘‘order’’: order(list), ‘‘groundtruth’’: groundtruth(str), ‘‘operations’’: number of the operations(int), }} === Please generate response for the given image that **strictly follow** the above requirments: